OpenStudy (faiqraees):

Is this method correct? \[\large\rm \frac{da}{dx}\frac{1}{a}=\frac{1}{\frac{da}{dx}a} \]

OpenStudy (faiqraees):

@mathmale@ganeshie8

hartnn (hartnn):

Derivative operator doesn't work that way. Except for limits, i don't think any operator works that way!

hartnn (hartnn):

is 'a' the function of x ?

OpenStudy (anonymous):

what is the question?

hartnn (hartnn):

if not, da/dx is just 0 ._.

hartnn (hartnn):

or did you mean \(\Large\rm \frac{d}{da}\frac{1}{a}=\frac{1}{\frac{d}{da}a}\) ??

hartnn (hartnn):

which is obviously not correct